OpenStudy (anonymous):

Is there a way to find the slope of the line determined by the following equation, without having to graph it all out on paper? 3x=4y-2

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

What formula did you use to get 3/4?

OpenStudy (anonymous):

Rewrite it as Y=(3/4)x+(2/4)

OpenStudy (anonymous):

Y=mx+c Where m is the gradient c is the y-axis intersect

OpenStudy (anonymous):

So basically you just isolate y instead of x?

OpenStudy (anonymous):

I still can't see how you got from 3x=4y-2 to Y=(3/4)x+(2/4) ...

OpenStudy (anonymous):

I've tried just isolating Y and did not come up with your formula...

OpenStudy (anonymous):

Add 2 to both sides 3x+2=4y Then dividie both sides by 4 (3/4)x+2/4=y Rewrite as Y=(3/4)x+1/2